OpenStudy (anonymous):

Evaluate the discriminant of 25x2 + 350x + 49 = 0. Tell how many solutions the equation has and whether the solutions are real or imaginar

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\]

hartnn (hartnn):

discriminant \(D = b^2-4ac = ...\)

hartnn (hartnn):

if D <0 roots are NOT real, imaginary if D>0 , roots are real. If D is a perfect square, roots are rational. If D = 0, roots are equal